If angles are supplementary, then one of the angles is an obtuse angle.
is this always true sometimes true or never true.?

Answer by umail08(87)

Supplementary angles: sum of angles is equal to 180 degrees
Obtuse angle: measure of angle is greater than 90 degrees and less than 180 degrees.
Therefore:
We can represent the supplement of an angle by the phrase:
(180-x) degrees
So we can test some angles to determine what is the supplement:
If: x=80 degrees, then (180-x) = (180-100) = 80 degrees One angle is obtuse
If: x=90 degrees, then (180-x) = (180-90) = 90 degrees Neither angle is obtuse
If: x = 100 degrees, then (180-x) = (180-100) = 80 degrees One angle is obtuse
So we see that there is one case where neither angle is obtuse …
Therefore the answer would be “sometimes true”